# Decomposing force fields as flows on graphs reconstructed from stochastic trajectories

### v1

@@ 1 Introduction
Concurrently, topological and graph-theoretic methods for the analysis of large and complex datasets have emerged as a powerful paradigm within data science [22, 23, 24, 25]. Such methods have also found applications in the analysis of time-series through techniques like attractor reconstruction [26] and, more recently, graph and topological signal processing [27, 28, 29, 30]. Amongst these methods, the Helmholtz-Hodge decomposition (HHD), which decomposes flows into gradients and circulations, has proved to be a valuable tool in the analysis of empirical signals evolving on discrete objects [31, 32, 33]. Surprisingly, an analogous decomposition to the HHD can also be applied to stationary Langevin processes, decomposing stochastic dynamics into reversible and irreversible parts [33, 18, 34]. In an attempt to bridge the HHD of an SDE with that of an associated flow on a graph, we construct a data-driven approach for approximating a Langevin process with a Markov chain constructed entirely from stochastic trajectories. Firstly, we use a maximum-likelihood estimator to reconstruct the transition intensities of a discrete-state, continuous-time Markov chain that approximates the underlying Langevin process. By taking an appropriate ratio of the rates, we define an edge-flow on a triangulation of phase-space and decompose this flow to gain insight into the irreversible and reversible forces driving the dynamics of the trajectories. Our approach is validated with numerical simulations of both solvable and nonlinear systems and then applied to the analysis of flickering red-blood cell (RBC) experiments and healthy and arrhythmic heartbeats. In particular, our approach is able to capture the difference in the prevalence of irreversible, circulating currents between the dynamics of healthy and adenosine triphosphate-depleted (ATP) RBCs [21, 35, 36] as well as the increased time-irreversibility of healthy heartbeats compared to arrhythmia [37, 38].
@@ 2.1 The various faces of the Helmholtz-Hodge decomposition
Figure 1: The Helmholtz-Hodge decomposition in various spaces. A. The HHD of a linear stable vector field $\bm{F}$ into a gradient component, $-\nabla\bm{\Phi}$, and a rotational component, $\bm{R}$. B. The HHD of a stationary Langevin process into its reversible component which climbs the stationary distribution (plotted behind) and its irreversible component which contours the distribution deterministically. The arrows in the final panel indicate the probability flux $\bm{J}$. C. The HHD of an edge-flow on a simplicial complex into gradient, curl and harmonic parts. Simplices are oriented naturally according to vertex labels. The gradient part sums to zero around closed loops. The curl part is only defined on triangles (example flow from Ref. [32]). D. The HHD of an edge-flow on a triangulation into gradient and curl parts. As the complex has no holes, there is no harmonic flow. Blank edges have zero flow.
@@
where $\bm{F}$ represents the deterministic, driving force of the system, known as the drift, whilst $\bm{\xi}\in\mathbb{R}^{n}$ are Gaussian fluctuations, known as the diffusion, with spatially-homogeneous correlations given by $\langle\bm{\xi}(t),\bm{\xi}(t^{\prime})\rangle=2\bm{D}\delta(t-t^{\prime})$ where $\bm{D}$ is the constant, positive definite, noise-covariance matrix. The density dynamics of the Langevin process are described the Fokker-Plank (FP) equation [2],
$\frac{\partial P(\mathbf{x},t)}{\partial t}=$ $-\sum_{i}\frac{\partial}{\partial x_{i}}\left[F_{i}P\right]+\sum_{i,j}\frac{\partial}{\partial x_{i}\partial x_{j}}\left[D_{ij}P\right].$
At stationarity, i.e. when $\partial P/\partial t=0$, we can decompose the drift as follows [34],
@@
where $\bm{F}_{\text{irr}}$ $\bm{F}_{\text{rev}}$ represent the time-irreversible and time-reversible components of the drift respectively, $P_{ss}$ is the stationary distribution and $\bm{J}_{ss}$ is the stationary probability flux (see App. A.1). More specifically, $\bm{F}_{\text{irr}}$ is odd whilst $\bm{F}_{\text{rev}}$ is even under time-reversal [34]. The detailed balance condition is satisfied if $\bm{J}_{ss}=0$, corresponding to an equilibrium stationary distribution with time-reversible dynamics, and broken if $\bm{J}_{ss}\neq 0$, corresponding to a nonequilibrium stationary distribution and time-irreversible dynamics [18]. Thus, information about the (ir)reversibility of the stochastic dynamics can be inferred by decomposing the drift field of the underlying stationary Langevin process. As shown in panel B of Fig. 1, the reversible component of the drift climbs the stationary distribution whilst the diffusion causes random fluctuations. The irreversible part of the drift drives deterministic rotation, contouring the stationary distribution.
In the case of constant isotropic noise, $\bm{D}=\epsilon\bm{I}$, the drift can be decomposed, as in the HHD, into the sum of divergence-free curl flux and a gradient flow,
@@ 2.2 Approximating force fields as flows on graphs
Figure 2: Reconstruction and decomposition of drift fields as flows on graphs. A. We begin with an ensemble of stochastic trajectories from the same system as a point-cloud on phase-space. B. We sub-sample the points using furthest point sampling to obtain a uniform density representation of phase-space. C. We discretise phase-space using a triangulation. Red points denote the center of each bin. D. We use maximum-likelihood estimation to reconstruct the transition intensities of a discrete-state Markov process with transitions between adjacent bins. E. Triangulating the midpoints of the bins, we obtain a simplicial complex representation of phase-space with transitions occurring over edges. F. Following the analysis in Sec. 2.1, we define the edge-flow by taking the logarithm of the square-root of the ratio of the transition intensities between connected bins. G. We perform the discrete HHD to obtain the curl and gradient components of the flow on the complex.
We attempt to leverage the correspondence of the HHD of the continuous Langevin and discrete Markov processes, by reconstructing a Markov chain approximation of the force field from an ensemble of stochastic trajectories and performing the discrete HHD. This will allows us to identify the curling and gradient components of the force field in a computationally simple manner. This is due to the relative ease with which the HHD of an edge-flow can be computed through its formulation as a least-squares optimisation problem (see App. A.1 and Ref. [31]).
@@
Next, we obtain the dual triangulation of the original grid by using the midpoints of the bins as vertices and performing the Delauney triangulation again. As a result, each vertex corresponds to a discrete state/bin in the Markov process and adjacent bins/reachable states are now connected by edges. Following the formulation of the HHD for Markov processes defined in Sec. 2.1, the flow on the edges of this second triangulation is defined to be $X_{ij}=\log\sqrt{q_{ij}/q_{ji}}$ where we assume the natural orientation $(i,j)$ if $i<j$. This flow can be decomposed with the HHD into its irreversible, curling component and its reversible, gradient component.
@@ 3.2 Real-world examples
In order to reconstruct a 2-dimensional trajectory in phase-space from a scalar measurement, we use a 2-d time-delay embedding with a single lag, $\tau$, chosen for each experiment by the mean of the first zero-crossing of the auto-correlation, thus we obtain a 2-d trajectory $(x(t),x(t+\tau))$ [58] (see App. A.6.1 for details). Panels B and D of Fig. 4 show examples of reconstructed trajectories for healthy and passive RBCs and healthy and arrhythmic heartbeats. Unlike the example processes in Sec. 3.1, we cannot assume the diffusion is isotropic. As described in Sec. 2.2.1, we first sub-sample and triangulate the trajectory in the reconstructed phase-space and use Eq. 10 to estimate the diffusion matrix. Next, we change coordinates, $\mathbf{y}=\bm{D}^{1/2}\mathbf{x}$, resample and re-triangulate to then perform the analysis assuming isotropic diffusion. In order to compare if the statistical structure of the trajectories in each condition is responsible for the variation in $P_{c}$, we perform two forms of significance testing. Firstly, we use ensemble bootstrapping [16, 17] with 100 bootstraps to estimate the uncertainty in $P_{c}$ in each case (see App. A.6.3). Additionally, we perform surrogate-testing with shuffled time-series where shuffling destroys the statistical structure whilst keeping the points in phase-space the same, thus generating a ‘null’ series [17] (see App. A.6.4). Furthermore, in App. A.6.2 we assess the validity of assuming spatially-homogeneous diffusion in these datasets. For further details on the datasets and the analysis see App. A.6.
@@ A.1.1 The HHD of a stationary Langevin process
The density dynamics of the Langevin process in Eq. 1 can be expressed by the FP equation, Eq. 2.1. In turn, this can be written as a probability conservation,
@@
which has zero divergence at stationarity. The drift at stationarity can be written as [18],
@@
i.e. as the sum of a divergence-free curl flux, $\bm{J}_{ss}/P_{ss}$, and the gradient of a scalar potential, $U=\log P_{ss}$, multiplied by the diffusion $\bm{D}$. Note that, whilst we assume spatially homogeneous diffusion, we include the term $\nabla\cdot\bm{D}(=0)$ for completeness. This term is also the correction factor between the Îto and Stratonovich formulations of spatially-dependent diffusions [66]. As a result, the stationary distribution and diffusive field give us the unique HHD of the drift.
@@ A.5 Variation of method parameters
Figure 5: Variation of method parameters. Here we show the effect of varying the number of sub-sampled points used to construct the grid and graph, the number of trajectories in an ensemble and the number of time-points in a trajectory, using the OU process with $\theta=4$ as an example. The estimate from the exact decomposition at $\theta=4$ is 0.652 (3.s.f).
@@ A.6.4 Surrogate-testing with shuffled data
A.6.5 Flickering experiments with red-blood cells

### v2

@@ 1 Introduction
Concurrently, topological and graph-theoretic methods for the analysis of large and complex datasets have emerged as a powerful paradigm within data science [22, 23, 24, 25]. Such methods have also found applications in the analysis of time-series through techniques like attractor reconstruction [26] and, more recently, graph and topological signal processing [27, 28, 29, 30]. Amongst these methods, the Helmholtz-Hodge decomposition (HHD), which decomposes flows into gradients and circulations, has proved to be a valuable tool in the analysis of empirical signals evolving on discrete objects [31, 32, 33]. Surprisingly, an analogous decomposition to the HHD can also be applied to stationary Langevin processes, decomposing stochastic dynamics into reversible and irreversible parts [33, 18, 34]. In an attempt to bridge the HHD of an SDE with that of an associated flow on a graph, we construct a data-driven approach for approximating a Langevin process with a Markov chain constructed entirely from stochastic trajectories. Firstly, we use a maximum-likelihood estimator to reconstruct the transition intensities of a discrete-state, continuous-time Markov chain that approximates the underlying Langevin process. By taking an appropriate ratio of the rates, we define an edge-flow on a triangulation of phase-space and decompose this flow to gain insight into the irreversible and reversible forces driving the dynamics of the trajectories. Our approach is validated with numerical simulations of both solvable and nonlinear systems and then applied to the analysis of flickering red-blood cell (RBC) experiments and healthy and arrhythmic heartbeats. In particular, our approach is able to capture the difference in the prevalence of irreversible, circulating currents between the dynamics of healthy and adenosine triphosphate-depleted (ATP) RBCs [21, 35, 36] as well as the increased time-irreversibility of healthy heartbeats compared to those with arrhythmia [37, 38].
@@ 2.1 The various faces of the Helmholtz-Hodge decomposition
Figure 1: The Helmholtz-Hodge decomposition in various spaces. A. The HHD of a linear stable vector field $\bm{F}$ into a gradient component, $-\nabla\bm{\Phi}$, and a rotational component, $\bm{R}$. B. The HHD of the drift field of a stationary Langevin process into its reversible component, the portion of the drift that maintains the stationary distribution (plotted behind - black) as constant despite the noisy fluctuations, and its irreversible component, which drive deterministic rotation around the stationary distribution. The first panel shows a trajectory from the stationary OU process, whilst the second and third show the same trajectory decomposed into its reversible and irreversible dynamics respectively. The arrows in the final panel indicate the stationary probability flux $\bm{J}_{ss}$. C. The HHD of an edge-flow on a simplicial complex into gradient, curl and harmonic parts. Simplices are oriented naturally according to vertex labels. The gradient part sums to zero around closed loops. The curl part is only defined on triangles (example flow from Ref. [32]). D. The HHD of an edge-flow on a triangulation into gradient and curl parts. As the complex has no holes, there is no harmonic flow. Blank edges have zero flow.
@@
where $\bm{F}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is a vector field on $\mathbb{R}^{N}$ that represents the time-constant, deterministic, driving force of the system, known as the drift, whilst $\bm{\xi}\in\mathbb{R}^{n}$ are noise in the form of Gaussian fluctuations, known as the diffusion, with spatially-homogeneous correlations given by $\langle\bm{\xi}(t),\bm{\xi}(t^{\prime})\rangle=2\bm{D}\delta(t-t^{\prime})$ where $\bm{D}\in\mathbb{R}^{N}\times\mathbb{R}^{N}$ is the constant, positive definite, noise-covariance matrix. Alternatively, one can consider the probability density of the process where $P(\mathbf{x},t)$ represents the probability of the process being in state $\mathbf{x}$ at time $t$. The density evolves according to the Fokker-Planck (FP) equation [2],
$\frac{\partial P(\mathbf{x},t)}{\partial t}=$ $-\sum_{i}\frac{\partial}{\partial x_{i}}\left[F_{i}P\right]+\sum_{i,j}\frac{\partial}{\partial x_{i}\partial x_{j}}\left[D_{ij}P\right].$ (2)
This can be written as a conservation of probability law,
$\frac{\partial P}{\partial t}+\nabla\cdot\bm{J}$ $=0,$ (3)
[…]
@@
where $\bm{F}_{\text{irr}}$ $\bm{F}_{\text{rev}}$ represent the time-irreversible and time-reversible components of the drift respectively, $\nabla$ is the gradient of a scalar potential (see App. A.1), $P_{ss}(\mathbf{x})$ is the stationary distribution, the solution to Eq. 2 when $\partial P/\partial t=0$, and $\bm{J}_{ss}(\mathbf{x})$ is the stationary probability flux, which satisfies $\nabla\cdot\bm{J}=0$. More specifically, $\bm{F}_{\text{irr}}$ is odd whilst $\bm{F}_{\text{rev}}$ is even under time-reversal [34]. The detailed balance condition is satisfied if $\bm{J}_{ss}\equiv\bm{0}$, corresponding to an equilibrium stationary distribution with time-reversible dynamics, and broken if $\bm{J}_{ss}\not\equiv\bm{0}$, corresponding to a nonequilibrium stationary distribution and time-irreversible dynamics [18]. Thus, information about the (ir)reversibility of the stochastic dynamics can be inferred by decomposing the drift field of the underlying stationary Langevin process. As shown in panel B of Fig. 1, the reversible component of the drift climbs the stationary distribution whilst the diffusion causes random fluctuations. The irreversible part of the drift drives deterministic rotation, contouring the stationary distribution. In the case of constant isotropic noise, $\bm{D}=\epsilon\bm{I}$, the drift can be written as the sum of a divergence-free curl flux ($\nabla\cdot\bm{J}_{ss}/P_{ss}=0$) and the gradient of a scalar potential,
@@ 2.2 Approximating force fields as flows on graphs
Figure 2: Reconstruction and decomposition of drift fields as flows on graphs. A. We begin with an ensemble of stochastic trajectories from the same system as a point-cloud on phase-space, illustrated here with two trajectories. B. We sub-sample the points using furthest point sampling to obtain a uniform density representation of phase-space. C. We discretise phase-space using a triangulation. Red points denote the center of each bin. D. We use maximum-likelihood estimation to reconstruct the transition intensities of a discrete-state Markov process with transitions between adjacent bins. E. Triangulating the midpoints of the bins, we obtain a simplicial complex representation of phase-space with transitions occurring over edges. F. Following the analysis in Sec. 2.1, we define the edge-flow by taking the logarithm of the square-root of the ratio of the transition intensities between connected bins. G. We perform the discrete HHD to obtain the curl and gradient components of the flow on the complex. The edge-flow colourbar refers to panels F and G.
Our approach attempts to approximate an unknown, underlying continuous Langevin process with a discrete Markov process that is reconstructed from an ensemble of stochastic trajectories. Representing the Markov process as a flow on a triangulation, we can leverage the correspondence between the HHD of the continuous and discrete-state processes to perform the HHD in discrete-space thus decomposing the process into its reversible and irreversible parts. This will allows us to approximate the curling and gradient components of the continuous drift-field with a discrete counterpart that can be calculated in a computationally simple manner. This is due to the relative ease with which the HHD of an edge-flow can be computed through its formulation as a least-squares optimisation problem (see App. A.1 and Ref. [31]).
@@
Next, we obtain the dual triangulation of the original grid by using the midpoints of the bins as vertices and performing the Delauney triangulation again. As a result, each vertex corresponds to a discrete state/bin in the Markov process and adjacent bins/reachable states are now connected by edges. The flow on the edges of this second triangulation is defined to be $X_{ij}=\log\sqrt{q_{ij}/q_{ji}}$ where we assume the natural orientation $(i,j)$ if $i<j$. This flow is the discrete counter-part to the drift field as its curling component represents irreversible flow and its gradient component represents reversible flow. More specifically, the Markov process obeys detailed balance iff this flow is conservative [33].
@@ 3.2 Real-world examples
In order to reconstruct a 2-dimensional trajectory in phase-space from a scalar measurement, we use a 2-d time-delay embedding with a single lag, $\tau$, chosen for each experiment by the mean of the first zero-crossing of the auto-correlation, thus we obtain a 2-d trajectory $(x(t),x(t+\tau))$ [58] (see App. A.6.1 for details and App. A.6.5 for the effect of varying this parameter.). Panels B and D of Fig. 4 show examples of reconstructed trajectories for healthy and passive RBCs and healthy and arrhythmic heartbeats. Unlike the example processes in Sec. 3.1, we cannot assume the diffusion is isotropic. As described in Sec. 2.2.1, we first sub-sample and triangulate the trajectory in the reconstructed phase-space and use Eq. 12 to estimate the diffusion matrix. Next, we change coordinates, $\mathbf{y}=\bm{D}^{1/2}\mathbf{x}$, resample and re-triangulate to then perform the analysis assuming isotropic diffusion. In order to compare if the statistical structure of the trajectories in each condition is responsible for the variation in $P_{c}$, we perform two forms of significance testing. Firstly, we use ensemble bootstrapping [16, 17] with 100 bootstraps to estimate the uncertainty in $P_{c}$ in each case (see App. A.6.3). Additionally, we perform surrogate-testing with shuffled time-series where shuffling destroys the statistical structure whilst keeping the points in phase-space the same, thus generating a ‘null’ series [17] (see App. A.6.4). Furthermore, in App. A.6.2 we assess the validity of assuming spatially-homogeneous diffusion in these datasets. For further details on the datasets and the analysis see App. A.6.
@@
4 Related work and other approaches
The analysis of nonequilibrium dynamics in biological systems has received much recent interest [18, 6], in particular in data-driven methods for analysing real-world trajectories [16, 8, 1, 59]. One typical approach is to infer approximate drift and diffusion terms directly from data [1, 13, 10, 9, 14], however only one such method ‘Stochastic Force Inference’ [8] allows for the natural HHD of the inferred drift as well as an estimate of the entropy production rate. Typically, methods aim to approximate a continuous field with a discrete approximation, but do not formulate a discrete-state stochastic process, as we do here. ‘Probability flux analysis’ [16] is another approach, designed for biological data, that attempts to approximate probability currents in a coarse-grained discrete state-space. This approach approximates the continuous process with this discrete-state counterpart, by inferring the probability flux between system states, yet it stops short of decomposing the process into its reversible and irreversible parts or measuring the overall level of nonequilibrium. More generally, many methods have been developed to analyse nonequilibrium flows by utilising the link between nonequilibrium dynamics and the irreversibility of stationary trajectories [60, 61] with applications in biology [19] and neuroscience [62]. Similarly to the quantity $P_{c}$, such methods can quantify the overall prevalence of nonequilibrium flows in stochastic trajectories, but yield no insight into the governing dynamics of the process, unlike the reconstruction of a discrete-state process.
The discrete HHD has found many applications in the analysis of hierarchical and cyclical structures in graphs and tournaments [63, 31] with some applications to dynamics and signal processing [32, 28, 29, 30]. To this point, the discrete HHD has not been applied to Markov processes to gain insight into their thermodynamics with the exception of Ref. [33]. Nevertheless, connections can be made between the HHD and classical ‘cycle decomposition’ approaches that are well-studied in the thermodynamics of graphs and Markov processes [64, 65, 66]. Similarly, results in the convergence of the discrete and continuous HHD are limited to simple domains and dynamics [33].
The approximation of vector fields with discrete counter-parts is a central challenge in numerical analysis, such as finite element calculus [67]. Such approaches typically involve discretisation of continuous domains for the numerical solution of deterministic equations and have little relation to our approach. One more closely related approach is ‘MARBLE’ [48] which aims to learn the manifold structure of phase-space, which is approximated as a graph, and infer the vector field directly from recordings of neural signals - though the dynamics are assumed to be deterministic thus the vector field is not formulated as a discrete-state Markov process.
@@ A.1.1 The HHD of a stationary Langevin process
As detailed in Eq. 1, the Langevin process is given by,
$\dot{\mathbf{x}}(t)$ $=\bm{F}(\mathbf{x}(t))+\bm{\xi}(t),$ (14)
where $\bm{F}$ is the drift field, whilst $\bm{\xi}$ are Gaussian fluctuations with spatially-homogeneous correlations given by $\langle\bm{\xi}(t),\bm{\xi}(t^{\prime})\rangle=2\bm{D}\delta(t-t^{\prime})$ with $\bm{D}$ the constant, positive definite, noise-covariance matrix. As shown in Eq. 2 and Eq. 3, the density dynamics can be written as a conservation of probability law,
@@
[…]
$\nabla\cdot\bm{A}(\mathbf{x})=\sum_{i}\partial\bm{A}/\partial x_{i},$ (17)
where $\mathbf{x}=(x_{1},...,x_{N})$. When defined on a tensor field, $\bm{D}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}\times\mathbb{R}^{N}$, it returns a vector in $\mathbb{R}^{N}$ whose $i$-th entry is given by,
$(\nabla\cdot\bm{D}(\mathbf{x}))_{i}$ $=\sum_{j}\frac{\partial D_{ij}(\mathbf{x})}{\partial x_{j}}.$ (18)
At stationarity, i.e. $\partial P/\partial t=0$, the density of the process remains constant whilst the trajectories still evolve in time. The probability flux at stationarity, $\bm{J}_{ss}$, is divergence-free, $\nabla\cdot\bm{J}_{ss}=0$. The drift at stationarity can be written as [18],
@@
where $\nabla$ represents the gradient operator, which takes a scalar potential, $U:\mathbb{R}^{N}\rightarrow\mathbb{R}$ and returns a vector field,
$\nabla U(\mathbf{x})$ $=\sum_{i}\frac{\partial U}{\partial x_{i}}\mathbf{e}_{i},$ (20)
where $\mathbf{e}_{i}$ is the $i-$th unit vector i.e. $(\mathbf{e}_{i})_{j}=\delta_{ij}$. Eq. 19 decomposes the drift field into the sum of a curl flux, $\bm{J}_{ss}/P_{ss}$, which is divergence-free, and the gradient of the scalar potential, $U=\log P_{ss}$, multiplied by the diffusion $\bm{D}$. Note that, whilst we assume spatially homogeneous diffusion, we include the term $\nabla\cdot\bm{D}(=0)$ for completeness. This term is also the correction factor between the Îto and Stratonovich formulations of spatially-dependent diffusions [74]. As a result, the stationary distribution and diffusive field give us the unique HHD of the drift. Moreover, in the case of a isotropic diffusion, $\bm{D}=\epsilon\bm{I},$ the reversible dynamics are exactly the gradient of the scalar potential $\log\epsilon P_{ss}$, and the drift is given by,
$\bm{F}=\bm{J}_{ss}/P_{ss}+\nabla\log\epsilon P_{ss}.$ (21)
As a result, the detailed balance condition, $\bm{J}_{ss}\equiv\bm{0}$, which guarantees reversible dynamics and an equilibrium stationary distribution, is met if and only if the drift field is conservative, i.e. there exists a scalar potential $V$ such that $\nabla V=\bm{F}$. This condition can be used to define a discrete-analog of the drift for the discrete-state Markov process and discrete gradient operator.
@@ A.5 Variation of method parameters
Figure 5: Variation of method parameters. Here we show the effect of varying the number of sub-sampled points used to construct the grid and graph, the number of trajectories in an ensemble and the number of time-points in a trajectory, using the OU process with $\theta=4$ as an example. The estimate from the exact decomposition at $\theta=4$ is 0.652 (3.s.f), as shown by the red dashed line.
@@ A.6.4 Surrogate-testing with shuffled data
A.6.5 Variation of the time-delay parameter
When analysing the real-world data, we used a single lag for each experiment that was chosen to be the mean first-zero-crossing of the autocorrelation function, which is a conventional choice [58]. For the RBC experiments we obtain $\tau=998,(T=10,000)$, whilst for the heartbeats we obtain $\tau=45,(T=188)$. In Fig. 8, we consider the effect of varying this parameter. For the RBCs, we consider lags of 250, 500, 750, 1000 and 1250. For the heartbeats, we consider lags of 15, 30, 45 and 60. For the RBCs, panel A shows that variation in the lag parameter shows no significant change in the distribution of $P_{c}$ over 100 boot-strapped samples, whilst the relative relationship between healthy and passive/fixed cells is kept the same. In panel B, we can see that variation in the lag parameter can cause greater variation in $P_{c}$ for the heartbeat dynamics. This is because the heartbeat data has a relatively short-time series which is made even shorter by an increase in the lag. This can increase the error in the measurements and lead to greater variation. Nevertheless, the qualitative relationship between normal and arrhythmic heartbeats appears to hold. This analysis suggests that our approach is not very sensitive to changes in the lag parameter provided the lag is relatively small in comparison with the length of the time-series.
Figure 8: Variation of the lag parameter, $\tau$, in the delay embedding. A. Varying the lag parameter in the RBC experiments shows little to no change suggesting that the phase-space dynamics are comparable for a range of different lags. B. In the case of the heartbeat dynamics, the time-series are much shorter as a result, varying the lag parameter causes variation in the quantity, but the relative relationship between normal and arrhythmic heartbeats is preserved.
A.6.6 Classifying time-series with standard methods
Figure 9: Classifying recordings with a neural network: We attempt to classify the time-series data considered here with a neural network trained on the summary statistics of the data. We find that for both the RBCs and the heartbeats, the neural network achieves almost perfect classification after very little training. This is due to clear differences in the summary statistics that allow for easy classification.
Whilst we have shown that our method can differentiate between recordings from biological systems in different conditions, our main goal was to gain insight into the underlying reversible and irreversible dynamics directly from stochastic trajectories. Our approach allows for insight into the thermodynamics and nonequilibrium physics of the biological processes, and goes beyond a classification tool [18]. Measuring $P_{c}$, the proportion of the flow in the curl space, is a first application that allows us to validate that our approach does capture irreversible flows by confirming results from the literature [37, 38]. We note that differentiating between conditions using the time-series data considered here is not a challenging task in itself, as we will show in this section. Beyond the prevalence of irreversible flows, the underlying time-series have significant differences in their statistics, that allow for easy classification. To illustrate this, we consider a neural-network classifier and show that it achieves (almost) perfect classification of recordings when trained on simple time-series summary statistics. For each of recording in the dataset, we compute a battery of trajectory features in the form of summary statistics. In particular, we consider the mean, median, standard deviation (Std), minimum value (Min), maximum value (Max), sum, skew, kurtosis and last value of each signal. For the RBCs, recordings are either healthy or passive, whilst for heartbeats, recordings are either normal or arrhythmic. We then train a neural-network on the summary statistics of a subset of the recordings with the goal of classifying each recording into its respective category. Our network has 2 hidden layers with 32 and 16 ‘ReLu’ units respectively and an output layer with sigmoidal activation. We use a binary cross-entropy loss function and the Adam optimiser. We train the network for 20 epochs with a validation split of 0.2 and a batch size of 50. In addition, we scale all recordings between [0,1] before training and use class-weighting to account for the different densities of the classes in the dataset. As shown in Fig. 9, we find that the neural network has (almost) perfect accuracy after only a few epochs of training. On the test set, we find that it classified both RBC and heartbeat recordings with 100% accuracy. Fig. 10 shows the distribution of the summary statistics, after data normalisation, for both healthy and passive cells. A subset of the statistics, namely the standard deviation, Min, Max and kurtosis are significantly different between healthy and passive cells, therefore, these features allow for simple classification of trajectories. Similarly, Fig. 11 shows the corresponding summary statistics for healthy and arrhythmic heartbeats, where the mean, median, standard deviation, sum, skew and kurtosis are all found to be significantly different.
Our approach is also able to distinguish between nonequilibrium dynamics such as those found in healthy RBCs, and the equilibrium dynamics of surrogate data generated by shuffling recordings in time. All the summary statistics considered in the previous experiment are invariant under shuffling of the data in time, which suggests that such trajectory features are insufficient for such differentiation. However, one can select summary statistics that depend on the temporal ordering, such as auto-correlation, to distinguish true and shuffled recordings. In Fig. 12, we show that the auto-correlation of data from Healthy 1, the RBC, is significantly higher than a surrogate generated by shuffling the data in time. This allows for the easy distinguishing of true dynamics from shuffled counterparts based on simple summary statistics.
Figure 10: Distribution of summary statistics for normalised red-blood cell data: We consider 10 summary statistics as features to input into the neural network. Here we show that some statistics, including standard deviation, Min, Max and kurtosis, are significantly different between healthy and passive cells, making for easy classification of the recordings. Significance is denotes as (ns) if $p>0.05$, (*) if $p<0.05$, (**) if $p<0.01$, (***) if $p<0.001$ and (****) if $p<0.0001$.
Figure 11: Distribution of summary statistics for normalised heartbeat data data: We consider 10 summary statistics as features to input into the neural network. Here we show that some statistics, including mean, median, standard deviation, sum, skew, kurtosis, and Last Value are significantly different between healthy and arrhythmic heartbeats, making for easy classification of the recordings. Significance is denotes as (ns) if $p>0.05$, (*) if $p<0.05$, (**) if $p<0.01$, (***) if $p<0.001$ and (****) if $p<0.0001$.
Figure 12: Distribution of auto-correlation for healthy RBC data and shuffled surrogate data: One can distinguish between true dynamics and shuffled surrogate data using temporally-dependent summary statistics. Here we show that the auto-correlation of data from the RBC Healthy 1 is significantly higher than its shuffled surrogate. Significance is denotes as (ns) if $p>0.05$, (*) if $p<0.05$, (**) if $p<0.01$, (***) if $p<0.001$ and (****) if $p<0.0001$.
A.6.7 Flickering experiments with red-blood cells